# A Task-Centric Theory for Iterative Self-Improvement with Easy-to-Hard Curricula

### v1

@@ Abstract
Iterative self-improvement fine-tunes an autoregressive large language model (LLM) on reward-verified outputs generated by the LLM itself. In contrast to the empirical success of self-improvement, the theoretical foundation of this generative, iterative procedure in a practical, finite-sample setting remains limited. We make progress toward this goal by modeling each round of self-improvement as maximum-likelihood fine-tuning on a reward-filtered distribution and deriving finite-sample guarantees for the expected reward. Our analysis reveals an explicit feedback loop where better models accept more data per iteration, supporting sustained self-improvement while explaining eventual saturation of such improvement. Adopting a task-centric view by considering reasoning tasks with multiple difficulty levels, we further prove quantifiable conditions on model initialization, task difficulty, and sample budget where easy-to-hard curricula provably achieve better guarantees than training on fixed mixtures of tasks. Our analyses are validated via Monte-Carlo simulations and controlled experiments on graph-based reasoning tasks.
@@ 1 Introduction
  • Empirical validation on graph-based reasoning tasks. We validate the analysis for both questions via controlled experiments on graph-based reasoning tasks (Section˜6) and complement these experiments with Monte-Carlo simulations that directly visualize the evolution of the expected reward lower bound across self-improvement iterations (Section˜5).
@@ 2 Related Work
While LLM self-improvement has shown strong empirical success, another line of work seeks to understand its underlying mechanisms from a theoretical perspective. Huang et al. [15] formalizes self-improvement as a consequence of a sharpening mechanism, which encourages the model to place larger probability mass on higher-quality sequences. From a different perspective, Mohri et al. [21] studies self-improvement through the lens of coherence. Sun et al. [27] models the training dynamics of self-improvement via the solver-verifier gap. In contrast to these works, our theoretical modeling of self-improvement is more closely aligned with mathematical reasoning settings, i.e., binary and verifiable rewards together with reject sampling for data collection. More importantly, our analysis covers multi-step iterative self-improvement and further incorporates its interaction with easy-to-hard curricula.
@@ 6 Experiment
6.1 Task and Training Setup
@@
Model and training.
We use Llama-3.2-1B-Instruct as our base LLM [12]. For each choice of the number of nodes $N$, expected out-degree $\bar{d}$, and target distance $l$, we generate a large collection of distinct graphs $\mathcal{G}$ together with vertex pairs $(v_{s},v_{t})$ whose shortest path length equals $l$. Each instance is rendered into natural language using a unified prompt template, forming a single shared sample pool. Across experiments, to obtain initialization models with diverse expected rewards $V_{p_{0}}(\hat{\theta}_{0})$, we warm up the base LLM by finetuning it on different subsets of the sample pool with varied finetuning hyperparameters, and draw task sets of varying difficulty from a disjoint portion of the pool. We then follow the procedure described in Section 3 and the easy-to-hard/baseline setup in Section 5.1 to run iterative self-improvement. More details of dataset construction, warm-up, and self-improvement finetuning are provided in Appendix D.
@@ 6.2 Experimental Results
Figure 3: Iterative self-improvement. Panel (a) shows the self-improvement trajectories of a fixed $\hat{\theta}_{0}$ across tasks with different initial Pass@1 accuracies; hollow markers and faded line segments indicate model collapse (Pass@1=0=0 for at least one target distance $l$). Panel (b) shows the performance under different question budgets $n$, with $\hat{\theta}_{0}$ and the initial Pass@1 fixed. Panel (c) shows the performance under different per-question answer budgets $m$, with $\hat{\theta}_{0}$ and the initial Pass@1 fixed.
In this section, we empirically validate that the main conclusions in Sections 4 and 5 are supported by the observed iterative self-improvement behavior on the shortest path task.
@@
Figure 3(a) shows the iterative self-improvement performance under tasks of varying difficulty, where we construct task sets such that the initial Pass@1 accuracy (corresponding to $V_{p_{0}}(\hat{\theta}_{0})$) ranges from $6\%$ to $78\%$. When the task is overly easy, we observe a decreasing trend in the test Pass@1 across iterations, whereas when the task is overly hard, the Pass@1 becomes unstable and may collapse. These trends are consistent with the analysis in Remark 4.2, which predicts that effective iterative self-improvement is only guaranteed to occur in a moderate difficulty regime. We also note that the improvement in Pass@1 often slows down after $t=2$ and the curves begin to plateau at values clearly below $1$, which is consistent with Remark 4.2.
Moreover, Figure 3(b)-(c) demonstrate that increasing either the question budget $n$ or the per-question answer budget $m$ consistently improves self-improvement performance. This aligns with the finite-sample interpretation in Section 4.1. It is worth mentioning that, despite minor differences in self-improvement setups, a range of empirical studies echo our findings on a broader set of real-world benchmarks that self-improvement tends to favor a moderate task difficulty regime [25], benefits from larger budgets ($n$ [25, 30] and $m$ [36, 2, 34]), and often exhibits a clear saturation limit [26].
@@
Figure 4: Iterative self-improvement with an easy-to-hard curriculum. Panel (a) fixes $\Delta=0.04$ and $\hat{\theta}_{0}$, and shows for different initial Pass@1 accuracies, the final Pass@1 gap between easy-to-hard and the baseline (i.e., $V_{p_{0}}(\hat{\theta}^{\mathrm{E2H}}_{L})-V_{p_{0}}(\hat{\theta}^{\mathrm{B}}_{L})$) as a function of the adjacent task difficulty ratio (captured by $\beta^{\prime}$). Panel (b) fixes $\Delta=0.04$ and $\beta^{\prime}=0.25$, and shows for different initial Pass@1 accuracies (spanning $35\%$$55\%$), how the final Pass@1 gap varies with the question budget $n$. The solid line reports the mean gap across 15 initializations and the shaded region indicates $\pm 1$ standard error; the gray bars (right axis) show the number of initializations with a positive gap at each $n$.
@@ 7 Conclusion
In this work, we developed a task-centric framework for understanding iterative self-improvement pipelines. For a single task, our finite-sample analysis characterizes key factors (e.g., task difficulty and sampling budget) that determine when (multi-step) self-improvement happens, and it also exposes intrinsic limits on what can be achieved through self-improvement alone. Beyond single-task training, we provide theoretical guidance for structuring iteration-wise curricula across difficulty levels and identify regimes where easy-to-hard scheduling yields a stronger lower bound guarantee than fixed-mixture training. By analyzing the feasibility and improvement conditions, we highlight the role of appropriate adjacent task difficulty ratios and a critical sample size near which the provably advantageous initialization region for easy-to-hard undergoes a sharp phase transition. Finally, our predictions are further supported by synthetic shortest path experiments, where we observe trends consistent with our theory.
@@ D.  Implementation Details and Training Setup
D.1  Graph Generation and Sample Pool
@@
D.2  Warm-up and Self-Improvement Finetuning Details
@@
LoRA dropout $0.05$
Table 2: Self-improvement finetuning hyperparameters used at each iteration.
@@
Next, for self-improvement, given an initialization model $\hat{\theta}_{0}$, we select an appropriate training set of size $nL$ together with a held-out test set such that the empirical Pass@1 accuracy of $\hat{\theta}_{0}$ matches the target value $V_{p_{0}}(\hat{\theta}_{0})$, enabling control of the initialization performance. We fix the test set across all $L$ iterations, with a test size of $1000$. To mitigate the effect of class imbalance, we enforce that, for both the per-iteration training set and the test set, the number of questions in each distance class $l\in\{-1,1,2,3\}$ is approximately balanced (up to a small tolerance); moreover, within each class, we also match (up to a small tolerance) the number of questions that are initially answered correctly by $\hat{\theta}_{0}$. For experiments comparing easy-to-hard against the baseline, we require a different form of control. Specifically, the training set of size $nL$ should admit two different partitions: (i) a baseline partition whose $L$ subsets (each containing $n$ questions for one iteration) have the approximately same (up to a small tolerance) initial Pass@1, matching $V_{p_{0}}(\hat{\theta}_{0})$; and (ii) an easy-to-hard partition whose initial Pass@1 values across iterations satisfy Assumption 5.1. we impose these constraints jointly and solve the resulting data selection problem using a CP-SAT solver.

### v2

@@ Abstract
Iterative self-improvement fine-tunes an autoregressive large language model (LLM) on reward-verified outputs generated by the LLM itself. In contrast to the empirical success of self-improvement, the theoretical foundation of this generative, iterative procedure in a practical, finite-sample setting remains limited. We make progress toward this goal by modeling each round of self-improvement as maximum-likelihood fine-tuning on a reward-filtered distribution and deriving finite-sample guarantees for the expected reward. Our analysis reveals an explicit feedback loop where better models accept more data per iteration, supporting sustained self-improvement while explaining eventual saturation of such improvement. Adopting a task-centric view by considering reasoning tasks with multiple difficulty levels, we further prove quantifiable conditions on model initialization, task difficulty, and sample budget where easy-to-hard curricula provably achieve better guarantees than training on fixed mixtures of tasks. Our analyses are validated through Monte-Carlo simulations and experiments spanning a synthetic graph-based reasoning task and multiple standard mathematical reasoning benchmarks.
@@ 1 Introduction
  • Empirical validation across diverse mathematical reasoning tasks. We validate the analysis for both questions through experiments on a synthetic graph-based reasoning task and multiple standard mathematical reasoning benchmarks, including GSM8K [3] and DeepMind Mathematics [25] (Section˜6), and complement these experiments with Monte-Carlo simulations that directly visualize the evolution of the expected reward lower bound across self-improvement iterations (Section˜5).
@@ 2 Related Work
While LLM self-improvement has shown strong empirical success, another line of work seeks to understand its underlying mechanisms from a theoretical perspective. Huang et al. [16] formalizes self-improvement as a consequence of a sharpening mechanism, which encourages the model to place larger probability mass on higher-quality sequences. From a different perspective, Mohri et al. [22] studies self-improvement through the lens of coherence. Yang et al. [38] investigates the impact of budget allocation policies across self-improvement iterations. Sun et al. [29] models the training dynamics of self-improvement via the solver-verifier gap. In contrast to these works, our theoretical modeling of self-improvement is more closely aligned with mathematical reasoning settings, i.e., binary and verifiable rewards together with reject sampling for data collection. More importantly, our analysis characterizes multi-step iterative self-improvement and further incorporates its interaction with easy-to-hard curricula.
@@ 6 Experiment
In this section, we empirically validate the main theoretical predictions developed in Sections 4 and 5 across diverse mathematical reasoning tasks. We begin with a synthetic shortest path task, which enables precise control over key variables in our theory. We then show that the resulting trends are also consistent with those observed on different standard mathematical reasoning benchmarks.
6.1 Experiments on Synthetic Tasks
6.1.1 Experimental Setup
@@
Datasets and model training setup.
Across experiments, we construct synthetic datasets from a sample pool consisting of a large collection of distinct graphs $\mathcal{G}$ together with vertex pairs $(v_{s},v_{t})$, spanning different choices of the number of nodes $N$, expected out-degree $\bar{d}$, and target distance $l$. This synthetic shortest path task enables direct control of several key variables, including the initialization expected reward $V_{p_{0}}(\hat{\theta}_{0})$, sample budgets $n$ and $m$, and difficulty ratios between adjacent tasks. We use Llama-3.2-1B-Instruct as our base LLM [13]. We follow the procedure described in Section 3 and the easy-to-hard/baseline setup in Section 5.1 to run iterative self-improvement. More details of dataset construction, parameter control, and self-improvement finetuning are provided in Appendix D.1.
@@ 6.1.2 Experimental Results
Figure 3: Iterative self-improvement results on the synthetic shortest path task. Panel (a) shows the self-improvement trajectories of a fixed $\hat{\theta}_{0}$ across tasks with different initial Pass@1 accuracies; hollow markers and faded line segments indicate model collapse (Pass@1=0=0 for at least one target distance $l$). Panel (b) shows the performance under different question budgets $n$, with $\hat{\theta}_{0}$ and the initial Pass@1 fixed. Panel (c) shows the performance under different per-question answer budgets $m$, with $\hat{\theta}_{0}$ and the initial Pass@1 fixed.
@@
Figure 3(a) shows the iterative self-improvement performance under tasks of varying difficulty, where we construct task sets such that the initial Pass@1 accuracy (corresponding to $V_{p_{0}}(\hat{\theta}_{0})$) ranges from $6\%$ to $78\%$. When the task is overly easy, we observe a decreasing trend in the test Pass@1 across iterations, whereas when the task is overly hard, the Pass@1 becomes unstable and may collapse. These trends are consistent with the analysis in Remark 4.2, which predicts that effective iterative self-improvement is only guaranteed to occur in a moderate difficulty regime. We also note that the improvement in Pass@1 often slows down after $t=2$ and the curves begin to plateau at values clearly below $1$, which is consistent with Remark 4.2. Moreover, Figure 3(b)-(c) demonstrate that increasing either the question budget $n$ or the per-question answer budget $m$ consistently improves self-improvement performance. This aligns with the finite-sample interpretation in Section 4.1.
@@
Figure 4: Iterative self-improvement with easy-to-hard curriculum on the synthetic shortest path task. Panel (a) fixes $\Delta=0.04$ and $\hat{\theta}_{0}$, and shows for different initial Pass@1 accuracies, the final Pass@1 gap between easy-to-hard and the baseline (i.e., $V_{p_{0}}(\hat{\theta}^{\mathrm{E2H}}_{L})-V_{p_{0}}(\hat{\theta}^{\mathrm{B}}_{L})$) as a function of the adjacent task difficulty ratio (captured by $\beta^{\prime}$). Panel (b) fixes $\Delta=0.04$ and $\beta^{\prime}=0.25$, and shows for different initial Pass@1 accuracies (spanning $35\%$$55\%$), how the final Pass@1 gap varies with the question budget $n$. The solid line reports the mean gap across 15 initializations and the shaded region indicates $\pm 1$ standard error; the gray bars (right axis) show the number of initializations with a positive gap at each $n$.
@@
6.2 Experiments on Standard Mathematical Reasoning Benchmarks
6.2.1 Experimental Setup
Datasets and model training setup.
We further validate our theoretical analysis on standard mathematical reasoning benchmarks, including GSM8K [3] and DeepMind Mathematics [25]. GSM8K contains 8.5K grade-school-level math word problems. DeepMind Mathematics consists of mathematical reasoning problems spanning multiple categories (e.g., algebra, arithmetic, and numbers) with each category further containing a large collection of fine-grained modules. Notably, DeepMind Mathematics comes with a native easy/medium/hard difficulty partition, which naturally supports our modeling of the easy-to-hard curriculum. Unless otherwise specified, we use Qwen3-8B [36] as the base model throughout Section 6.2. Additional base models considered in this section include Llama-3.2-1B-Instruct, Llama-3.2-3B-Instruct [13], Qwen2.5-3B-Instruct, Qwen2.5-7B-Instruct [37], and Qwen3-32B [36]. As in Section 6.1, we follow the setup in Section 3 and Section 5.1 to run iterative self-improvement, and use Pass@1 accuracy as the evaluation metric. More details on dataset splits and self-improvement finetuning are provided in Appendix D.2.
Model Iteration $t$ Absolute Improvement
0 1 2 3
Llama-3.2-1B-Instruct 2.96 3.56 3.14 2.68 -0.28
Llama-3.2-3B-Instruct 6.07 6.95 5.97 5.61 -0.46
Qwen2.5-3B-Instruct 10.39 11.78 13.04 13.62 -3.23
Qwen2.5-7B-Instruct 18.12 19.16 20.17 20.19 -2.07
Qwen3-8B 27.52 28.66 29.75 30.67 -3.15
Qwen3-32B 38.59 41.47 43.09 43.77 -5.18
Table 1: Self-improvement trajectories of different base models on GSM8K. Iteration $t$ reports the model’s Pass@1 accuracy before iteration $t$ (equivalently, after iteration t1t{-}1). The last column reports the absolute improvement between the final Pass@1 and the initial value. All values are percentages.
Model Iteration $t$ Absolute Improvement
0 1 2 3
Llama-3.2-1B-Instruct 18.89 35.16 36.42 36.09 -17.20
Llama-3.2-3B-Instruct 37.00 46.04 51.00 54.44 -17.44
Qwen2.5-3B-Instruct 55.89 61.29 67.31 70.42 -14.53
Qwen2.5-7B-Instruct 74.00 76.98 81.44 82.98 - 8.98
Qwen3-8B 85.11 86.65 90.89 92.78 - 7.67
Qwen3-32B 97.89 95.67 96.24 96.69 -1.20
Table 2: Self-improvement trajectories of different base models on DeepMind Mathematics Dataset. Iteration $t$ reports the model’s accuracy before iteration $t$ (equivalently, after iteration t1t{-}1). The last column reports the absolute improvement between the final value and the initial value. All values are percentages.
$n$ Iteration $t$ Absolute Improvement
0 1 2 3
600 27.52 21.59 22.30 23.46 -4.06
1200 27.52 26.76 27.32 28.08 -0.56
1800 27.52 28.40 29.13 30.25 -2.73
2400 27.52 28.66 29.75 30.67 -3.15
(a) Different question budgets $n$.
$m$ Iteration $t$ Absolute Improvement
0 1 2 3
1 27.52 28.66 29.75 30.67 -3.15
4 27.52 29.02 31.13 31.51 -3.99
16 27.52 30.42 31.55 32.86 -5.34
(b) Different per-question answer budgets $m$.
Table 3: Self-improvement trajectories of Qwen3-8B under different budgets on GSM8K. Iteration $t$ reports the model’s Pass@1 accuracy before iteration $t$ (equivalently, after iteration t1t{-}1). The last column reports the absolute improvement between the final Pass@1 and the initial value. All values are percentages.
6.2.2 Experimental Results
Iterative self-improvement.
Table 1 and Table 2 report the self-improvement trajectories on GSM8K and on the easy split of the nine DeepMind Mathematics modules with the highest initial Qwen3-8B accuracy, respectively. Across different base models, these trajectories clearly show that when a task is too easy or too hard relative to the model, self-improvement tends to be limited or even deteriorate, whereas moderate task difficulty is more conducive to sustained iterative self-improvement. Moreover, as the number of iterations increases, the incremental improvement from self-improvement tends to decrease. Table 3 further shows how the degree of self-improvement changes with the question budget $n$ and the per-question answer budget $m$. The results indicate that larger values of $n$ and $m$ are both more favorable to self-improvement.
Overall, these findings are consistent with the synthetic task results in Section 6.1.2, and provide empirical support for Remark 4.2, Remark 4.2, and the finite-sample analysis in Section 4.1. It is worth noting that, despite minor differences in self-improvement setups, a range of empirical studies further echo our findings on a broader set of real-world benchmarks: self-improvement tends to favor a moderate task difficulty regime [27], benefits from larger budgets ($n$ [27, 32] and $m$ [41, 2, 39]), and often exhibits a saturation limit [28].
Category Module $\rho$
0.0 0.2 0.4 0.6 0.8 1.0
algebra linear_2d -0.00 -0.04±0.27 -0.84±0.11 -2.52±0.54 -1.96±0.27 -2.04±0.25
arithmetic add_sub_multiple -0.00 -2.76±2.23 -7.68±2.97 -11.60±3.17 -17.08±2.12 -10.64±3.64
div -0.00 -0.16±0.45 -0.42±0.50 -0.98±0.61 -0.88±0.49 -0.54±0.50
mixed -0.00 -1.06±1.10 -1.64±1.36 -3.40±1.20 -4.62±0.51 -3.20±1.54
mul_div_multiple -0.00 -2.60±1.62 -4.76±2.41 -9.56±2.20 -12.32±2.09 -10.48±2.33
nearest_integer_root -0.00 -2.38±2.80 -2.50±1.09 -2.72±2.14 -4.44±3.14 -6.64±2.55
numbers div_remainder -0.00 -2.80±0.90 -2.16±0.72 -3.78±0.42 -3.80±0.88 -2.90±0.67
gcd -0.00 -0.22±0.96 -0.32±0.62 -0.78±0.74 -0.04±0.91 -0.44±1.11
lcm -0.00 -0.12±0.60 -0.26±0.65 -1.04±0.65 -0.96±0.86 -2.66±1.04
place_value -0.00 -0.02±0.73 -0.28±0.73 -0.02±0.54 -0.44±0.78 -3.12±1.52
Table 4: Final Pass@1 gap between easy-to-hard and the baseline on the DeepMind Mathematics Dataset using Qwen3-8B. Difficulty follows the dataset’s native easy/medium/hard split. $\rho$ interpolates between fixed-mixture training in every round ($\rho{=}0$, baseline) and a fully staged curriculum ($\rho{=}1$: iteration 0 easy, iteration 1 medium, iteration 2 hard). We run experiments on all modules in the algebra, arithmetic, and numbers categories. To focus on modules for which curriculum effects are meaningfully measurable, we exclude modules whose canonical difficulty split yields $\beta^{\prime}<0.1$, or whose gap variation across $\rho$ is smaller than 1%. Underlined entries denote, for each module, the maximum gap across $\rho$. All values are percentages. Standard errors are shown as subscripts.
Figure 5: The left panel plots the estimate of $\beta^{\prime}$ as a function of $\rho$, and the right panel plots the estimate of $\beta$ as a function of $\rho$, for all modules reported in Table 4. Here, the estimates of $\beta^{\prime}$ and $\beta$ are computed from the initialization model using the form in Assumption 5.1.
Iterative self-improvement with easy-to-hard curriculum.
Beyond the synthetic experiments, here we further validate our analysis using the dataset’s native difficulty partition. Concretely, for the DeepMind Mathematics dataset, we vary the difficulty ratios between adjacent tasks through a parameter $\rho$. For the $L=3$ iteration self-improvement setting, let $\mathbf{W}=(w_{ij})\in\mathbb{R}^{3\times 3}$, where $w_{ij}$ denotes the sampling weight assigned at iteration i1i-1 to the $j$-th difficulty level, with $j=1,2,3$ corresponding to the easy, medium, and hard splits, respectively. Let $\mathbf{U}$ be the matrix whose entries are all $1/3$, and let $\mathbf{I}$ be the identity matrix. We define
$\mathbf{W}=(1-\rho)\mathbf{U}+\rho\mathbf{I},\qquad\rho\in[0,1].$
Thus, when $\rho=0$, we have $\mathbf{W}=\mathbf{U}$, which corresponds to the baseline where each iteration samples uniformly from all three difficulty levels; when $\rho=1$, we have $\mathbf{W}=\mathbf{I}$, which corresponds to a fully staged easy-to-hard curriculum in which each iteration samples exclusively from one difficulty level. As $\rho$ increases from 0 to $1$, the difficulty ratios between adjacent tasks become larger.
Figure 5 shows that the estimates of $\beta$ and $\beta^{\prime}$ both increase clearly with $\rho$, suggesting that these $\beta$ and $\beta^{\prime}$ provide effective proxies for the dataset’s native notion of difficulty separation. Table 4 further reports, for modules in the algebra, arithmetic, and numbers categories, the final Pass@1 gap between easy-to-hard and the baseline as a function of $\rho$. Across modules, as $\rho$ increases from 0 to $1$, the final Pass@1 gap typically exhibits a trend of first increasing and then decreasing, with the largest gain attained at an interior value of $\rho$. These results support Remark 5.2, which predicts that moderate difficulty ratios between adjacent tasks are the most favorable for easy-to-hard scheduling to realize its advantage within the iterative self-improvement framework.
@@ 7 Conclusion
In this work, we developed a task-centric framework for understanding iterative self-improvement pipelines. For a single task, our finite-sample analysis characterizes key factors (e.g., task difficulty and sampling budget) that determine when (multi-step) self-improvement happens, and it also exposes intrinsic limits on what can be achieved through self-improvement alone. Beyond single-task training, we provide theoretical guidance for structuring iteration-wise curricula across difficulty levels and identify regimes where easy-to-hard scheduling yields a stronger lower bound guarantee than fixed-mixture training. By analyzing the feasibility and improvement conditions, we highlight the role of appropriate adjacent task difficulty ratios and a critical sample size near which the provably advantageous initialization region for easy-to-hard undergoes a sharp phase transition. Finally, our predictions are further supported by experiments on a synthetic shortest path task and multiple standard mathematical reasoning benchmarks, where we observe trends consistent with our theory.
@@ D.  Implementation Details and Training Setup
D.1  Experiments on Synthetic Tasks
Graph Generation and Sample Pool.
@@ D.1  Experiments on Synthetic Tasks
Warm-up and Self-Improvement Finetuning Details.
@@
LoRA dropout $5\times 10^{-2}$
Table 6: Self-improvement finetuning hyperparameters used at each iteration on synthetic tasks.
@@
Next, for self-improvement, given an initialization model $\hat{\theta}_{0}$, we select an appropriate training set of size $nL$ together with a held-out test set such that the empirical Pass@1 accuracy of $\hat{\theta}_{0}$ matches the target value $V_{p_{0}}(\hat{\theta}_{0})$, enabling control of the initialization performance. We fix the test set across all $L$ iterations, with a test size of $1000$. To mitigate the effect of class imbalance, we enforce that, for both the per-iteration training set and the test set, the number of questions in each distance class $l\in\{-1,1,2,3\}$ is balanced; moreover, within each class, we also match the number of questions that are initially answered correctly by $\hat{\theta}_{0}$. For experiments comparing easy-to-hard against the baseline, we require a different form of control. Specifically, the training set of size $nL$ should admit two different partitions: (i) a baseline partition whose $L$ subsets (each containing $n$ questions for one iteration) have the same (up to a small tolerance) initial Pass@1, matching $V_{p_{0}}(\hat{\theta}_{0})$; and (ii) an easy-to-hard partition whose initial Pass@1 values across iterations satisfy Assumption 5.1. we impose these constraints jointly and solve the resulting data selection problem using a CP-SAT solver.
@@
D.2  Experiments on Standard Mathematical Reasoning Benchmarks
Dataset split details.
GSM8K [3] is released with 7473 training examples and 1319 test examples. Each GSM8K question is rendered into a natural language prompt using a unified template. The prompt template is:
Solve the following grade-school math word problem.
Give only the final answer as an integer. Do not show steps.
[…]
Answer:
Throughout our GSM8K experiments, we fix the official test split for evaluation. For training, we fix the number of self-improvement iterations to $L=3$, and partition the training data into three rounds such that the model’s initial performance on each round matches its initial performance on the full training set. Except in the experiment that explicitly varies the question budget $n$, we set $n=2400$.
DeepMind Mathematics [25] contains multiple categories of mathematical reasoning problems, each further divided into fine-grained modules, with a native easy/medium/hard difficulty partition. To facilitate evaluation, we filter out all modules whose answers are not single numeric values. For the remaining modules, each question is rendered into a natural language prompt using the following unified template:
Solve the following school-level math problem.
Give only the final answer as a single integer, decimal, or fraction. Do not show steps.
[…]
Answer:
We construct a relatively simple task that yields the high initial model performance reported in Table 2 by selecting data from the easy splits of the nine modules on which Qwen3-8B achieves the highest accuracy. These modules are algebra_linear_1d, arithmetic_add_or_sub, arithmetic_div, arithmetic_mul, arithmetic_nearest_integer_root, measurement_conversion, numbers_gcd, numbers_place_value, and numbers_round_number. Our training and test sets contain samples drawn uniformly from these modules. We then partition the training data into three iterations such that the model’s initial performance on each round matches (up to a small tolerance) its initial performance on the selected training set. For Table 2, we set $n=400$. For the iterative self-improvement with easy-to-hard curriculum experiments, we consider the three largest categories, namely algebra, arithmetic, and numbers, and run experiments on each module in these categories. The data construction procedure is described in Section 6.2.2. For Table 4, we set $n=1000$.
Self-Improvement Finetuning Details.
Similar to Appendix D.1, for both GSM8K and DeepMind Mathematics, we perform iterative self-improvement exactly following the setup in Sections 3 and 5.1. Across all experiments on standard mathematical reasoning benchmarks, we use the same self-improvement finetuning hyperparameters summarized in Table 8.
Hyperparameter Value
Learning rate $2\times 10^{-4}$
Weight decay $1\times 10^{-4}$
Batch size 88
LoRA rank $16$
LoRA scaling $32$
LoRA dropout $5\times 10^{-2}$
Table 8: Self-improvement finetuning hyperparameters used at each iteration on standard mathematical reasoning benchmarks.